QUESTION 8

An object is dropped from rest, after how much time (in unit of second) will the final velocity be 21.1 m/s?
Answer #1

Given that object is dropped => follows gravitation

Dropped from rest => Initial velocity,u = 0

From kinematics, V = u+at

V =gt => t = 21.1/9.81 = 2.15 sec